OpenStudy (ebydaqueen19):

HELP HELP HELP!!!! One fruit salad recipe calls for 1/2 cup of sugar. Another recipe calls for 2 tablespoons of sugar. Since 1 tbsp is 1/16 cup, how much more sugar does the first recipe require?

OpenStudy (anonymous):

then 2 tbsp=1/8 cup 1/2-1/8=3/8 so the first recipe requires 3/8 cup more sugar

OpenStudy (ebydaqueen19):

Thank you :)

OpenStudy (anonymous):

You're welcome!